OpenStudy (anonymous):

The Fiedler family has up to $130,000 to invest. They decide that they want to have at least $40,000 invested in stable bonds yielding 5.5% and that no more than $60,000 should be invested in more volatile bonds yielding 11%. How much should they invest in each type of bond to maximize income if the amount in the stable bond should not exceed the amount in the more volatile bond? What is the maximum income?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

this is all i have so far... x = money invested in stable bonds (5.5%) y = money invested in volatile bonds (11%) total z = 0.055x + 0.11y x + y < 130,000 0.055x + y > 40,000 0.11y + x < 60,000

OpenStudy (anonymous):

This is a math problem, it's precalculus, it's a method to find the minimize the cost and maximize profit of something.

OpenStudy (anonymous):

we need to follow the following the procedures: 1Choose the unknowns. 2Write the objective function. 3Write the constraints as a system of inequalities. 4 Find the set of feasible solutions that graphically represent the constraints. 5 Calculate the coordinates of the vertices from the compound of feasible solutions. 6 Calculate the value of the objective function at each of the vertices to determine which of them has the maximum or minimum values. It must be taken into account the possible non-existence of a solution if the compound is not bounded.

OpenStudy (anonymous):

Corner (x, y) Objective Function "z = 86x + 130y" (x, y) z = 86(0) + 130(0) (0, 100) z = 86(0) + 130(100) = $ 13086 (85.56, 0) z = 86(85.56) + 130(0) = $ 7358.16 (0, 110) z = 86(0) + 130(110) = $ 14300 (35, 65) z = 86(35) + 130(65) = $ 11460 (100, 0) z = 86(100) + 130(0) = $ 8600 <----- Minimum cost. This is an example of another linear programing problem, after we set up the inequality, we will begins to find vertices and substitute x and y axis variables. It is easier to solve rather than yahoo's method.

OpenStudy (anonymous):

I have been looking around the internet, i found the correct answer for this question is: $60,000 in the stable bonds and $60,000 in the volatile bonds; maximum income $9900

OpenStudy (anonymous):

I hope you guys can correct my constraints of inequality to see if it is correct or not, i can do the rest. Hope you guys help me. Thank! x = money invested in stable bonds (5.5%) y = money invested in volatile bonds (11%) total z = 0.055x + 0.11y x + y < 130,000 0.055x + y > 40,000 0.11y + x < 60,000

ganeshie8 (ganeshie8):

I am getting below constraints : z = 0.055x + 0.11y x + y <= 130,000 x >= 40,000 y <= 60,000 0.055x <= 0.11y

OpenStudy (anonymous):

Okay, please hold on, i'm reviewing your constraints, hopefully i will get some ideas. =)

ganeshie8 (ganeshie8):

lets keep them even more simpler : z = 0.055x + 0.11y x + y <= 130,000 x >= 40,000 y <= 60,000 x <= y

ganeshie8 (ganeshie8):

they should work ^^

ganeshie8 (ganeshie8):

also, changing the units to thousands simplifies calculations : z = 0.055x + 0.11y x + y <= 130 x >= 40 y <= 60 x <= y

OpenStudy (anonymous):

Okay, i understand your constraints, i will work on the rest of the problem, and then I'll share with you my graphs and answers... Thanks! I gave you a medal as well. ^ ^

ganeshie8 (ganeshie8):

good luck ! you should get the MAXIMUM value for z at x=60, y=60 : http://www.wolframalpha.com/input/?i=maximize+z+%3D+0.055x+%2B+0.11y%2C+x+%2B+y+%3C%3D+130%2Cx+%3E%3D+40%2C+y+%3C%3D++60%2C+x+%3C%3D+y

OpenStudy (anonymous):

Yah, thank you so much for sharing your ideas, and as well as the website you just posted. I see this tool will help me a lot in my studying. I'm sure your answer is correct because i have been looking around the internet for the most reliable answer. And it is "$60,000 in the stable bonds and $60,000 in the volatile bonds; maximum income $9900" ^^

ganeshie8 (ganeshie8):

sounds good :) wolfram is your best friend when doing tough problems like these

OpenStudy (anonymous):

Thanks! for the medal..

OpenStudy (anonymous):

But after i graph this problem, How do i shade it? Which vertices (x,y) do i need to use to test?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

I know that and the intersection of (60,60) is where the Fiedler family should invest $60,000 in the stable bonds and $60,000 in the volatile bonds. But we must also use others (x,y) to test to prove that (60,60) is the most profitable.

ganeshie8 (ganeshie8):

yes, we need to test all the vertices of common region

ganeshie8 (ganeshie8):

graph the inequalities here : https://www.desmos.com/calculator

OpenStudy (anonymous):

I'm seeing 8 points in there, is it necessary to test out 8 points? or we should just pick the relevant one?

ganeshie8 (ganeshie8):

dont graph the line, graph the inequality...

ganeshie8 (ganeshie8):

i am getting only 3 points... wait a sec, il attach the graph

OpenStudy (anonymous):

Yeah, only 3 points he'll attach a graph

OpenStudy (anonymous):

Okay =( I'm waiting patiently.

ganeshie8 (ganeshie8):

that triangle is the common region satisfying all the constraints ^^

ganeshie8 (ganeshie8):

the maximum value occurs only at the vertices of that triangle

ganeshie8 (ganeshie8):

vertices of common region : (60, 60) (40, 60) (40, 40)

ganeshie8 (ganeshie8):

so you need to test the "z" at all those 3 points and see which point gives the maximum

OpenStudy (anonymous):

Yah, now i get it, your graph looks exactly like mine, how coincident, :P. Okay i will use those 3 points. But May i just shade the triangle instead everything else in different colors?

ganeshie8 (ganeshie8):

yes :) geogebra gives more visually pleasing picture... let me attach the graph from geogebra

OpenStudy (anonymous):

Okay, you have many secret tools =P.

OpenStudy (anonymous):

It seem like i need to download this software before using it. May i ask you if it is trustable?

ganeshie8 (ganeshie8):

100% trustable

ganeshie8 (ganeshie8):

there is an online version also : http://www.geogebra.org/webstart/geogebra.html

OpenStudy (anonymous):

Yah, I'm testing it now. =)